OpenStudy (hba):

The center of the circle with eqn. 4x^2+4y^2-8x+12y-25=0 is I am just conforming my answer...........

OpenStudy (anonymous):

No, I won't tell you..

OpenStudy (hba):

i did not ask

hartnn (hartnn):

posting a question means asking....

OpenStudy (hba):

Please see the ques carefully

OpenStudy (anonymous):

Where is the answer you want to confirm??

OpenStudy (hba):

(1,-3/2) < here you go :S

OpenStudy (anonymous):

If you have solved, then please put your answer in the question box..

OpenStudy (hba):

Ok :)

OpenStudy (anonymous):

It is looking bad to me..

OpenStudy (hba):

That means it is absolutely right :D

OpenStudy (anonymous):

If you think so then yes..

OpenStudy (hba):

radius of the circle with the eqn, x^2+y^2+12x-10y=0 is I got \[\sqrt{61}\]

OpenStudy (hba):

@waterineyes Right or wrong ?

OpenStudy (anonymous):

36 + 25 = ??

OpenStudy (hba):

61

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

Yes you are right radius is root(61)..